OpenStudy (banjoman):

A rectangular room is 1.9 times as long as it is wide, and its perimeter is 25 meters. Find the dimension of the room. will medal

OpenStudy (rcp031):

L=1.9W P=2L+2W=25, 2(1.9W)+2W=25, 5.8W=25, W=125/29=4.31 L=1.9*4.31=8.19